OpenStudy (precal):

what is the derivative of y=(tanx)/(1+secx)

OpenStudy (anonymous):

hint:multiply top/bottom by secx

OpenStudy (precal):

I got y' = (secx)/(1+secx) but wolfram alpha does not have that as a solution

OpenStudy (precal):

I can't multiply by secx for the top and bottom I have to use quotient rule

OpenStudy (anonymous):

sec(x) (tan^2(x)+sec^2(x)+sec(x))

OpenStudy (precal):

No Jinnie I don't think that is correct

OpenStudy (anonymous):

1/2 sec^2(x/2)

OpenStudy (anonymous):

hahaha i forgot it was a division problem

OpenStudy (precal):

ok how did you get that

OpenStudy (y2o2):

\[{-\sec(x) \tan(x)^2\over(\sec(x)+1)^2}+{\sin(x)^2\over(\cos(x)^2}+{\cos(x))+1\over(\sec(x)+1)}\]

OpenStudy (precal):

y2o2 what are you using to find y prime?

OpenStudy (anonymous):

it does work at walfarm, check this out: http://www.wolframalpha.com/input/?i=d%2F+dx+%28sinx%2Fcosx%29%2F%281%2Bsecx%29&cdf=1

OpenStudy (anonymous):

-ln|cosx+1|+c ?

OpenStudy (precal):

no iHelp that is not correct

OpenStudy (precal):

I used a pythagorean identity for tan^2t

OpenStudy (anonymous):

tanx/(secx+1)=sinx/cosx+1, let u=cosx+1, du=-sinxdx, so integral of -1/u, which is -ln|u| , subisitute u back in...

OpenStudy (precal):

this is not an integration problem, I am just looking for the derivative

OpenStudy (anonymous):

wow, srry misread the problem - my bad

OpenStudy (precal):

its ok, you were on a roll there

OpenStudy (anonymous):

just make sure, derivative of product is first*deri of second minus second*deri of first right?>

OpenStudy (anonymous):

y'=sec^2x(1+secx)-tan^2xsecx , is that right ?

OpenStudy (precal):

that is the product rule, I need to use the quotient rule.

OpenStudy (anonymous):

nvm. miss read problem again.. wow fml...

OpenStudy (precal):

I use lodehi-hidelo over lolo

OpenStudy (anonymous):

last try lol... y'=(sec^2x(1+secx)-tan(x-secxtanx))/(1+secx)^2,,

OpenStudy (precal):

yes I distributed tanx to secxtanx and got secxtan^2x then I used the pyth. identity to get rid of tan^2x

OpenStudy (precal):

I think I am right. I do believe wolfram alpha is not able to put all variations of a solution down. Programming issue.

OpenStudy (precal):

Thanks to everyone Good night

OpenStudy (anonymous):

=) solution page is not always right.

OpenStudy (anonymous):

night

OpenStudy (dumbcow):

yes you are correct...i also get dy/dx = sec/(1+sec) you can change it in terms of cos --> dy/dx = 1/(1+cos) and this is one of the forms given on wolframalpha

OpenStudy (precal):

when do you suggest I change it to cosx? In the beginning before I take the derivative or at the end?

OpenStudy (dumbcow):

it doesn't really matter...it depends on the function, in this case if you changed it before it would have made taking the derivative simpler

OpenStudy (precal):

ok I will go back and redo the problem. Thanks for your help :)